OpenStudy (anonymous):

How many real-number solutions does the equation have? -9x^2+10x+2=0

OpenStudy (anonymous):

use the quadratic equation \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

which ones would be a b and c?

OpenStudy (anonymous):

if your descriminant \[\sqrt{b ^{2}-4ac}\] is a positive value then you have a real value. If not the value is imaginary a,b,c correspond to the values of the form \[ax ^{2}+bx+c=0\]